OpenStudy (anonymous):

F of x equals the integral from 1 to x of the natural logarithm of t squared. Use your calculator to find F ′(3).

OpenStudy (anonymous):

https://gyazo.com/3d3327245c9230958f68094e39259155

OpenStudy (anonymous):

@jim_thompson5910 I selected that by mistake, I am not sure what to do here.

OpenStudy (anonymous):

@ganeshie8 how'd you get that? I got 13.2 plugging in 3

jimthompson5910 (jim_thompson5910):

By the Fundamental Theorem of Calculus (FTC) if \[\Large F(x) = \int_{a}^{x}f(t)dt\] then \[\Large F \ ' (x) = f(x)\]

ganeshie8 (ganeshie8):

** \(F(x) = \int\limits_{1}^x\ln (t^2)\,dt\) \(F'(x) = \ln(x^2)*(x)' = \ln(x^2)\) plugin \(x=3\)

ganeshie8 (ganeshie8):

A bit more generally : \[\dfrac{d}{d\color{red}{x}}\int\limits_a^{g(\color{red}{x})} f(t) \,dt = f(g(x))*g'(x)\]

OpenStudy (anonymous):

Ok, got it. Thanks to you both.

ganeshie8 (ganeshie8):

np